The King Food Corp. currently has no debt in its capital structure. The beta of its capital is 1.5. The King Food Corp’s free cash flow is expected to equal £30 million next year. This cash flow is expected to grow at 2% per year for the foreseeable future. King Food Corp. is considering changing its capital structure by issuing debt and using the proceeds to buy back stock. It will do so in such a way that it will have a 50% debt-to-equity ratio (D/E=50%) after the change, and it will maintain this debt-equity ratio forever. Assuming King Food Corp’s pretax cost of debt will be 5%. King Food Corp. faces a corporate tax rate 40%. Assuming that the CAPM holds, the risk-free rate is 3%, and the expected market index risk premium is 8%.


(a)  Calculate the WACC before the change in capital structure and after change in the capital structure, respectively.                                                            

                                   



1
Expert's answer
2022-03-04T12:17:55-0500

Solution

Given that

The beta of capital is     "=1.5"

Risk free rate            = 3%

Expected market index risk premium = 8%

Cash flow growth  = 2% pa

 

The cost of the equity before the change in capital structure is calculated as

CARM          = 3 + 1.5 (8 – 3)

= 3 + 7.5

= 10 .5

 

After the change, the

Debt to equity ratio D/E = 50% = 0.5

Corporate Tax rate = 40%


And the cost of the equity after the change in capital structure is calculated as

WACC         = 0.5 × 5 (1 – 0.4) + 0.5 × 10.5

                      = 0.5 × 3 + 5.25

                       = 6.75